you count 187 CFUS on a spread plate. The plate was prepared by spreading 0.2ml of a 1:100 dilution of a 1:50 dilution of a 1:10 dilution of the original sample. What is the concentration of the original sample. What is the concentration of the original culture in cells/ml
Answer #1

Given,

the number of colonies on the plate = 187

The volume of sample plated = 0.2 mL

Dilution used = 1 / 100 * 1 / 50 * 1 / 10 = 1 / 50000

Dilution here is calculated using the product of all the dilutions made. Dilution factor is the inverse of dilution. Here, 1 / 50000 means that the sample is diluted 50000 times.

The concentration of the original culture can be determined using the formula:

CFU / mL = No. of colonies * DF / Volume plated

where DF is the dilution factor.

incorporating each of the values in the above formula,

CFU / mL = 187 * 50000 / 0.2 = 9350000 / 0.2 = 46750000

The above value can be represented in scientific notation by using the first digit as the whole number and the rest as decimals and using the number of decimal digits as the number to the power ten. In this case, there are a total of 8 digits, so we are going to take the first as the whole number and take the remaining 7 as decimals and so it would be represented as :

4.675 * 107 CFU / mL

Now, each viable bacterial cell woulld form a CFU or Colony forming unit and so it would be the same as number of cells per mL, so the above answer can also be written in terms of cells / mL.

Therefore, the original concentration of the culture would be 4.675 * 107 cells / mL.